OpenStudy (anonymous):

which ordered pair is the solution of the equation? -7x-5y=42 a.) 0,-6 b.) -5,0 c.) 0, -5 d.) -6,0

OpenStudy (anonymous):

plug in each values for "x" and "y" in the left hand side and see if you get 42 (a) \(-7(0)-5(-6)=30\ne42\) (b) \(-7(-5)-5(0)=35\ne42\) (c) \(-7(0)-5(-5)=25\ne42\) (a) \(-7(6)-5(0)=42\)

OpenStudy (anonymous):

so, which pair satisfies the equation?